What force parallel to the inclined is required to push a 40 kg up on a 50o inclined plane if the coefficient of friction (static) is 0.3.


1
Expert's answer
2021-09-27T09:02:09-0400

"\vec F+m\vec g+\vec N+\vec{F_{fr}}=\vec 0,"

"Ox:F-F_{fr}-mg\sin\alpha=0,"

"Oy:N=mg\cos\alpha,"

"F_{fr}=\mu N,"

"F=mg(\mu\cos\alpha+\sin \alpha)=376~N."